254. Rudolph dies in 2018, having made taxable gifts of $3 million during his lifetime and having no taxable estate. An election is made on Rudolph’s estate tax return to permit Deborah, his wife to Rudolph's deceased spousal unused exclusion amount. As of Rudolph’s death, Deborah has made no taxable gifts. Therefore, Deborah's applicable exclusion amount, which she may use for lifetime gifts or for transfers at death is what amount?

A. $0B. $8,180,000C. $11,180,000D. $19,360,000
Answer #1

Answer:

Correct answer is:

D. $19,360,000

Explanation:

Given that:

An election is made on Rudolph’s estate tax return to permit Deborah, his wife to Rudolph's deceased spousal unused exclusion amount

Unused exclusion amount = 11,180,000 - 3000000 = $8,180,000

Deborah has made no taxable gifts. Hence she has her own life time exclusion amount of $11,180,000.

Deborah's applicable exclusion amount, which she may use for lifetime gifts= 11180000 + 8180000 =$19,360,000

Option D is correct.

Option A, B, C are incorrect as election is made on Rudolph’s estate tax return to permit Deborah to Rudolph's deceased spousal unused exclusion amount, Deborah's applicable exclusion = $19,360,000 as calculated above.